Suppose you are an astronaut and you have been stationed on a distant planet. You would like to find the acceleration due to the gravitational force on this planet so you devise an experiment. You throw a rock up in the air with an initial velocity of 9 m/s and use a stopwatch to record the time it takes to hit the ground. If it takes 5.2 s for the rock to return to the same location from which it was released, what is the acceleration due to gravity on the planet?
Answer #1

initial velocity u = 9 m/s

total time taken t = 5.2 s

time taken to reach the maximum height t' = t/2

t' = 2.6 s

use equation of motion

v = u + at;

0 = 9 + a x 2.6 (at the maximum height velocity is zero)

a = 3.462 m/s2   (acceleration due to gravity on the planet)